OpenStudy (anonymous):

Katie's history class has 24 students. If the ratio of boys to girls is 3:1, how many more boys are there than girls in the class?

OpenStudy (mathstudent55):

A ratio of 3:4 can be expressed as a fraction, \(\dfrac{3}{4} \) Let's look at some numbers that are in a ratio of 3:4 \(\dfrac{30}{40} \), \(\dfrac{6}{8} \), \(\dfrac{9}{12} \) Ok so far?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got the answer but thnx

OpenStudy (mathstudent55):

Each of those numbers is in a ratio of 3:4 because each one is written as a fraction that can be reduced to \(\dfrac{3}{4} \). For example, for \(\dfrac{30}{40}\), this is how you'd reduce it. \(\dfrac{30}{40} = \dfrac{3}{4} \times \dfrac{10}{10} = \dfrac{3}{4} \times \dfrac{\cancel{10}}{\cancel{10}} = \dfrac{3}{4}\)

OpenStudy (mathstudent55):

You can express the ratio \(\dfrac{3}{4}\) as \(\dfrac{3x}{4x} \). If numbers in a ratio add to a sum, you express it like this: 3x + 4x = sum In your case the ratio is 3:1, so you'd do 3x + x = 24. Solve for x. Then 3x is the number of boys and x is the number of girls. Then subtract.

OpenStudy (mathstudent55):

You're welcome.